# Differential pumping for kHz operation of a Laser Wakefield accelerator based on a continuously flowing Hydrogen gas jet

### v2

@@ III.1 Fluid dynamics simulations
In order to obtain more accurate estimates of the gas flow in the nozzle and the differential chamber, we have performed numerical simulations using the computational fluid dynamics software ANSYS Fluent. Simulations were performed using molecular hydrogen modeled by the Aungier-Redlich-Kwong real gas equation. 3D geometry was used. The results are shown in Figure 6 for $\mathrm{H_{2}}$ with a backing pressure $P_{back}=100\,$bar and a nozzle throat of $D^{*}=60\,\text{\textmu}$ µm. In the simulations, the boundary conditions at the entrance of the cones and the pumping outlet were taken to be $P=0$. The simulation results shown in Figure 6a confirm that the pressure remains at the mbar level in the differential chamber, away from the region where the high-density jet is present. Figure 6b shows a dramatic decrease of the density by three orders of magnitude in the gap region between the two cones. The density then steadily decreases along the cone which is favorable for coupling the intense laser into the high-density gas jet without prior interaction. Thus, the gas remains well-confined within the differential chamber, but there is still some stagnation observed about 1 cm before the gas jet, potentially impacting laser performance.
@@ III.2 Propagation of the laser pulse up to the gas jet
To evaluate whether the stagnating gas affects laser propagation, we have performed simulations of the propagation of the laser pulse from the parabola (at $z=-10\,$cm) to the jet (at $z=-200\,\text{\textmu}$ µm) using an optical propagation code. The simulations were conducted using three gases : $\mathrm{H}$, $\mathrm{He}$ and $\mathrm{N}$ in order to assess the potential impact of different gases on the laser pulse. The propagation of the laser through an initially neutral gas was modeled using the open-source library “Axiprop” Andriyash (2024). This library uses various optical propagators, depending on the specific conditions (gas or vacuum for instance) and concatenate them to handle more complex scenarios. For our study, the paraxial propagator was used for vacuum propagation from the parabola ($z=-10\,$cm) to the point where the gas density becomes non-negligible ($z=-10\,$mm), as it handles drastic changes in beam size Goodman (2005). For the region with gas ($z=-10\,$mm to $z=-200\,\text{\textmu}$ µm), a non-paraxial propagator was applied, following the method by Oubrerie and al. Oubrerie et al. (2022b). Both propagators were implemented using Hankel transforms to account for cylindrical symmetry Oubrerie et al. (2022b). To model the propagation of the laser in the gas/plasma, the solver that propagates the field within the plasma is similar to the one described by Couairon and al. Couairon et al. (2011). Ionization of the gas is simulated using the ADK model Ammosov, Delone, and Krainov (1986); Augst et al. (1989). In the gas region, grid sizes were set to $\Delta z=5.0\,\text{\textmu}$ µm for $\mathrm{He}$ and $\mathrm{N}$, $\Delta z=4\,\text{\textmu}$ µm for $\mathrm{H}$, and $\Delta r=0.64\,\text{\textmu}$ µm in all three gases.
@@
Figure 7 illustrates the evolution of the laser intensity ($I$) and waist ($w_{0}$) along its propagation through the density profiles of $\mathrm{N}$, $\mathrm{He}$, and $\mathrm{H}$ gases, with vacuum included for reference. Overall, the intensity and beam size remain nearly identical across vacuum and gas media, indicating that the effect of residual gas is largely negligible. However, at the entrance of the gas jet, where gas density inevitably rises, small differences become noticeable, causing slight perturbations in the laser pulse. Figure 7a shows a slight reduction in laser intensity when propagating through gas compared to vacuum. This reduction is most pronounced in $\mathrm{H}$ (reduction about 17%), and there is a smaller drop of roughly 8% in $\mathrm{N}$ and $\mathrm{He}$. Figure 7b highlights a slight increase in beam waist due to ionization-induced defocusing, with an increase of about 7% in $\mathrm{H}$, 5% in $\mathrm{N}$, and 4% in $\mathrm{He}$ relative to vacuum. Finally, Figure 7c shows that $\mathrm{H}$ fully ionizes early in the propagation, around $z=-4.4\,$mm, while $\mathrm{He}$ and $\mathrm{N}$ reach full ionization closer to $z=-700\,\text{\textmu}$ µm. These slight differences explain the laser pulse’s more noticeable degradation after propagating through $\mathrm{H}$, compared to $\mathrm{N}$ or $\mathrm{He}$.
@@
Also, the interaction of the pulse with all three gases shifts the spectrum towards the blue. Figure 8 shows the spectrum of the initial pulse and of the pulse at the end of the simulations, integrated over the radial axis, after propagating through vacuum, $\mathrm{N}$, $\mathrm{He}$, or $\mathrm{H}$. When propagating through the three gases, the spectrum shifts slightly to higher frequencies, with a maximum normalized frequency shift of $\omega_{0,\text{max}}/\omega_{0}=1.02$. The laser pulse behaves similarly in all three gases, with no significant differences regarding the final spectrum.
The stagnating gas has a minimal impact on laser propagation. While the interaction of the laser pulse with the gases results in a slight reduction in intensity, some increase in beam size, and a minor blueshift in the spectrum, these changes are negligible. The effects are slightly more pronounced in $\mathrm{H}$ due to its lower ionization threshold and the fact that it fully ionizes significantly earlier than $\mathrm{He}$ and $\mathrm{N}$.
@@ IV.1 Experimental set-up
The electron beam charge, spatial profile, and beam pointing are measured using a calibrated YAG screen imaged onto a 14 bits CCD camera. The electron energy spectrum is measured onto the same YAG screen, by inserting a motorized magnetic spectrometer in the beam, which relies on a set of permanent magnetic dipoles creating a magnetic field of 0.120.120.12\,0.12T on a 202020\,20mm length. Each image presented here was obtained by acquiring data over 111\,1ms (so each image corresponds to a single shot). For each plasma density studied, a series of 100 acquisitions is taken. Statistics over electron beam parameters are then obtained by averaging over the acquisition series, and the uncertainties represent the RMS deviation from the mean value.
@@ IV.2 Results
For each measurement, we optimized the laser pulse duration and focal position. We explored densities ranging from $n_{e}=7.4\times 10^{19}\,\mathrm{cm}^{-3}$ to $n_{e}=1.2\times 10^{20}\,\mathrm{cm}^{-3}$. Figure 9a shows a typical electron beam profile at the higher density of $n_{e}=1.2\times 10^{20}\,\mathrm{cm}^{-3}$. The beam exhibits a quasi-circular shape with minimal divergence, with very good stability shot to shot as shown in Table 1. Figure 9b shows the corresponding electron spectra for two different densities, averaged over 100 shots. The thickness of the lines corresponds to the RMS fluctuations in the dataset. Both spectra reveal low fluctuations, with well-defined energy peaks, confirming the consistent beam quality. The stability of the beam is further demonstrated in Figure 9c, where we present 100 consecutive single-shot acquisitions. The low variance in beam parameters such as charge, divergence, and energy indicates the high level of stability achieved with this setup. Table 1 summarizes the relative fluctuations of key beam parameters for the two plasma densities. Notably, the beam-pointing fluctuations remain below 2.3 mrad RMS, with charge fluctuations less than 7%, and energy fluctuations as low as 2% shot-to-shot, confirming excellent beam stability throughout the experiment.
@@ V Conclusion
Finally, we conducted laser wakefield acceleration experiments to test the efficacy of the differential pumping setup, yielding highly stable and well-defined electron beams. This accomplishment represents the first continuous operation at kHz repetition rates with hydrogen, thereby marking a significant advancement in kHz LWFA using light gases.
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement JRA PRISE no. 871124 Laserlab-Europe and IFAST under Grant Agreement No 101004730. This project was also funded by the Agence Nationale de la Recherche under Contract No. ANR-20-CE92-0043-01, and benefited from the support of Institut Pierre Lamoure via the chair “Accélération laser-plasma haute cadence”.

### v3

@@ III.1 Fluid dynamics simulations
In order to obtain more accurate estimates of the gas flow in the nozzle and the differential chamber, we have performed numerical simulations using the computational fluid dynamics software ANSYS Fluent. Simulations were performed using molecular hydrogen modeled by the Aungier-Redlich-Kwong real gas equation. 3D geometry was used. The results are shown in Figure 6 for $\mathrm{H_{2}}$ with a backing pressure $P_{back}=100\,$bar and a nozzle throat of $D^{*}=60\,\text{\textmu}$ µm. In the simulations, the boundary conditions at the entrance of the cones and the pumping outlet were taken to be $P=0$. The simulation results shown in Figure 6a confirm that the pressure remains at the mbar level in the differential chamber, away from the region where the high-density jet is present. Figure 6b shows a dramatic decrease of the density by three orders of magnitude in the gap region between the two cones. The density then steadily decreases along the cone which is favorable for coupling the intense laser into the high-density gas jet without prior interaction. Thus, the gas remains well-confined within the differential chamber, but there is still some stagnation observed about 1 cm before the gas jet, potentially impacting the laser pulse quality at the focus.
@@ III.2 Propagation of the laser pulse up to the gas jet
To evaluate whether the stagnating gas affects laser propagation, we performed simulations of the propagation of the laser pulse from the parabola (at $z=-10\,$cm) to the jet (at $z=-200\,\text{\textmu}$ µm) using an optical propagation code. The simulations were conducted using three gases: $\mathrm{H}$, $\mathrm{He}$ and $\mathrm{N}$ in order to assess the potential impact of different gases on the laser pulse. The propagation of the laser through an initially neutral gas was modeled using the open-source library “Axiprop” Andriyash (2024). This library uses various optical propagators, depending on the specific conditions (gas or vacuum for instance) and concatenate them to handle more complex scenarios. For our study, the paraxial propagator was used for vacuum propagation from the parabola ($z=-10\,$cm) to the point where the gas density becomes non-negligible ($z=-10\,$mm), as it handles drastic changes in beam size Goodman (2005). For the region with gas ($z=-10\,$mm to $z=-200\,\text{\textmu}$ µm), a non-paraxial propagator was applied, following the method by Oubrerie and al. Oubrerie et al. (2022b). Both propagators were implemented using Hankel transforms to account for cylindrical symmetry Oubrerie et al. (2022b). To model the propagation of the laser in the gas/plasma, the solver that propagates the field within the plasma is similar to the one described by Couairon and al. Couairon et al. (2011). Gas ionization is simulated using the ADK model Ammosov, Delone, and Krainov (1986); Augst et al. (1989). In the gas region, grid sizes were set to $\Delta z=5.0\,\text{\textmu}$ µm for $\mathrm{He}$ and $\mathrm{N}$, $\Delta z=4\,\text{\textmu}$ µm for $\mathrm{H}$, and $\Delta r=0.64\,\text{\textmu}$ µm in all three gases.
@@
Figure 7 illustrates the evolution of the laser intensity ($I$) and waist ($w_{0}$) along its propagation through the density profiles of $\mathrm{N_{2}}$, $\mathrm{He}$, and $\mathrm{H_{2}}$ gases, with vacuum included for reference. Overall, the laser intensity and beam size remain similar when propagating through vacuum and gas media. However, small differences in laser perturbations become apparent after propagation. Figure 7a shows a reduction in laser intensity when propagating through gas compared to vacuum. This reduction is most pronounced in $\mathrm{H_{2}}$ (reduction about 19% at $z=-200\,$µm), and there is a smaller drop of roughly 10% in $\mathrm{N_{2}}$ and $\mathrm{He}$. Figure 7b highlights a slight increase in beam waist due to ionization-induced defocusing, with an increase of about 8% in $\mathrm{H_{2}}$, and 4% in $\mathrm{N_{2}}$ and $\mathrm{He}$ relative to vacuum: along this propagation, $\mathrm{H_{2}}$ has a greater impact on the laser pulse than He or $\mathrm{N_{2}}$, though the overall perturbations remain limited.
@@
While this might initially seem counterintuitive, the key lies in the dynamics of the ionization process. The analysis tracks the evolution of the laser pulse from $z=-10\,$mm to $z=-200\,$µm before the center of the gas jet. Importantly, ionization occurs continuously along the laser propagation path, rather than at a single point. Additionally, the neutral gas density in simulations with $\mathrm{H_{2}}$ or $\mathrm{He}$ must be five times higher than that of $\mathrm{N_{2}}$ to maintain consistent electronic density at the focus, since these different molecules release different numbers of electrons upon ionization. As shown in Figure 7c, $\mathrm{H_{2}}$ molecules are fully ionized from $z=-6\,$mm onward, whereas $\mathrm{N_{2}}$ molecules are ionized only up to their first level there. As long as the laser ionizes only a single level of $\mathrm{N_{2}}$, it behaves similarly to when propagating in $\mathrm{H_{2}}$, but in a lower neutral gas density. On this interval, ionization has a greater impact in $\mathrm{H_{2}}$ due to its higher neutral gas density. This dynamics shifts around $z=-1\,$mm, where the laser fully ionizes $\mathrm{He}$ and $\mathrm{N_{2}}$ (up to its fifth level). Up to $z=-1\,$mm, the laser pulse experiences greater degradation in $\mathrm{H_{2}}$ compared to $\mathrm{He}$ or $\mathrm{N_{2}}$. Beyond this point, ionization-induced distortions of the laser pulse become more pronounced in $\mathrm{N_{2}}$, particularly in the main density peak Monzac et al. (2024), counterbalancing the earlier trends observed during the initial propagation.
Also, the interaction of the pulse with all three gases shifts the spectrum towards the blue. Figure 8 shows the spectrum of the initial pulse and of the pulse at the end of the simulations, integrated over the radial axis, after propagating through vacuum, $\mathrm{N_{2}}$, $\mathrm{He}$, or $\mathrm{H_{2}}$. When propagating through the three gases, the spectrum shifts slightly to higher frequencies, with a maximum normalized frequency shift of $\omega_{0,\text{max}}/\omega_{0}=1.02$. The laser pulse behaves similarly in all three gases, with no significant differences regarding the final spectrum.
The stagnating gas has a limited impact on laser propagation. While the interaction of the laser pulse with the gases results in a reduction in intensity, some increase in beam size, and a minor blueshift in the spectrum, these changes do not prevent a priori electron acceleration. The effects are slightly more pronounced in $\mathrm{H_{2}}$ due to its higher neutral gas density and to the fact that it fully ionizes significantly earlier than $\mathrm{He}$ and $\mathrm{N_{2}}$.
@@ IV.1 Experimental set-up
The electron beam charge, spatial profile, and beam pointing were measured using a calibrated YAG screen imaged onto a 14 bits CCD camera. The electron energy spectrum is measured onto the same YAG screen, by inserting a motorized magnetic spectrometer in the beam, which relies on a set of permanent magnetic dipoles creating a field of 0.120.120.12\,0.12T on a 202020\,20mm length. Each image presented here was obtained by acquiring data over 111\,1ms (so each image corresponds to a single shot). For each plasma density studied, a series of 100 acquisitions is taken. Statistics over electron beam parameters were then obtained by averaging over the acquisition series, and the uncertainties represent the RMS deviation from the mean value.
@@ IV.2 Results
For each measurement, we optimized the laser pulse duration and focal position by adjusting the jet position in 555\,5µm steps, while varying the laser pulse chirp in $1.6\,\mathrm{fs^{2}}$ steps. We explored densities ranging from $n_{e}=7.4\times 10^{19}\,\mathrm{cm}^{-3}$ to $n_{e}=1.2\times 10^{20}\,\mathrm{cm}^{-3}$. Figure 9a shows a typical electron beam profile at the higher density of $n_{e}=1.2\times 10^{20}\,\mathrm{cm}^{-3}$. The beam exhibits a quasi-circular shape with minimal divergence, as well as a remarkably good shot-to-shot stability, as shown in Table 1. Figure 9b shows the corresponding electron spectra for two different densities, averaged over 100 shots. The thickness of the lines corresponds to the RMS fluctuations in the dataset. Both spectra reveal low fluctuations, with well-defined energy peaks, confirming the consistent beam quality. The beam stability is further demonstrated in Figure 9c, where we present 100 consecutive single-shot acquisitions. The low variance in beam parameters such as charge, divergence, and energy indicates the high level of stability achieved with this setup. Table 1 summarizes the relative fluctuations of key beam parameters for the two plasma densities. Notably, the beam-pointing fluctuations remain below 2.3 mrad RMS, with charge fluctuations less than 7%, and energy fluctuations as low as 2% shot-to-shot, confirming excellent beam stability throughout the experiment.
@@ V Conclusion
Finally, we conducted laser wakefield acceleration experiments to test the efficacy of the differential pumping setup, yielding highly stable and well-defined electron beams. We hereby demonstrated the first continuous operation at kHz repetition rates with hydrogen, thereby marking a significant advancement in kHz LWFA using light gases.
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement JRA PRISE no. 871124 Laserlab-Europe and IFAST under Grant Agreement No 101004730. This project was also funded by the Agence Nationale de la Recherche under Contract No. ANR-20-CE92-0043-01, and benefited from the support of Institut Pierre Lamoure via the “Accélération laser-plasma haute cadence” fund.